ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 46, 47 Kraljevic, Hrvoje, University of Zagreb, Croatia Landau’s type inequalities for contraction cosine functions We consider Landau’s type inequalities for powers of infinitesimal generators of strongly continuous cosine functions of linear contractions on a Banach space. The constants are computed for powers less than or equal to 10. Reference: H. Kraljevic, On Landau’s type inequalities for infinitesimal generators, Mathematical Communications 1(1996), 111-120.